40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16794     Entry time: Thu Apr 21 11:31:35 2022
Author: JC 
Type: Update 
Category: VAC 
Subject: Gauges P3/P4 

[Jordan, JC]

It was brought to attention during yesterday's meeting that the pressures in the vacuum system were not equivalent althought the valve were open. So this morning, Jordan and I reviewed the pressure gauges P3 and P4. We attempted to recalibrate, but the gauges were unresponsive. Following this, we proceeded to connect new gauges on the outside to test for a calibration. The two gauges successfully calibrated at atmosperic pressure. We then proceeded to remove the old gauges and install the new ones. 

Attachment 1: IMG_0560.jpeg  2.537 MB  Uploaded Thu Apr 21 12:38:17 2022  | Hide | Hide all
IMG_0560.jpeg
Attachment 2: IMG_0561.jpeg  2.604 MB  Uploaded Thu Apr 21 12:38:19 2022  | Hide | Hide all
IMG_0561.jpeg
ELOG V3.1.3-